# Examples of robotic art Encountering Robotic Art: The Social, Material, and Temporal Processes of Creation with Machines

### v1

@@ 3.1. Participant Recruitment
In this study, we conducted in-depth, semi-structured interviews with nine artists working on robotic art. They are renowned professional artists who use or build physical robots for/their artistic work, namely robotic art. Designers and amateurs are out of this criterion. We identified potential participants by browsing individual and artistic organizations’ websites during the initial recruitment. Subsequently, we employed snowball sampling to recruit more artists who met our criteria. The recruitment ended when iterative data analysis reached saturation. The vast majority of the participants have had formal art training in drawing, painting, installation, media art, music, sculpture, or dance, and all of them incorporated robots into their particular forms. Some have technical backgrounds. Many have a doctoral degree in the fields of art or technology. All of them build their robots or at least modify out-of-shell robots. Table 1 shows more demographic information. We refrain from disclosing more specific information about the artists, e.g., art forms, and job titles, which reasonably increases the likelihood of identifying the individuals. The study is approved by the Institutional Review Board of the lead author’s institution. The participants provided oral consent to participate in this study voluntarily.
@@ 3.2. Semi-Structured Interviews
Before the interviews, the first author reviewed the media press on the artists’ websites, gaining background information to build rapport, develop interview questions, and inform analysis. He conducted the interviews via Zoom with video recording, which lasted an average of 67 minutes. The in-depth interview enabled us to look at the layers of meaning that artists relate with their work—unraveling feelings, attitudes, values, and motivations crucial for addressing our research question (Pugh, 2013). The initial interview questions are structured around the following topics: the workflow, motivation, rewards, and challenges of working on robotic art; the relationship between artists and robots; and the source of creativity. These inquiries were driven by our curiosity in making sense of the artists’ work practices. We aimed to understand their motivations for engaging with robotic art, to unpack both the explicit and implicit values embedded in their practices, and to explore broader topics around human-machine creation.
At the beginning of the interviews, the first author asked the participants to introduce themselves and their artwork. Then he asked about their inspiration source for their artwork and motivations for working on robotic art, steering the conversation toward the creativity and characteristics of robotic art. Following that, he solicited participants’ reflections and anecdotes on the opportunities and challenges of working on robotic art to understand the nuances of robotic art practices. He then guided the conversation to explore the relationships between artists and robots, focusing on their respective agencies and contributions to creative outcomes in art practices. Additionally, he posed questions regarding the artistic practices’ embodied/material engagement. Concluding the session, he asked the participants to elaborate on any aspects they wished to add to the discussion or to highlight anything missed. To comply with the semi-structured nature, he prompted follow-up questions throughout the interviews and remained open to new ideas introduced by the participants, allowing some emergent ideas to lead the discussion. After each interview, the participant received a $40 Amazon or Visa e-gift card if they chose to accept compensation.
@@ 3.3. Qualitative Analysis
Data collection and analysis proceeded in parallel and iterative ways throughout the study. We began the analysis shortly after conducting each interview. Emerging insights from the analysis led us to refine our interview protocol several times, focusing on the most salient story in the data. The analysis also informed theoretical sampling (Charmaz, 2006) at the later stage, purposefully recruiting participants more likely to provide missing components or potentially revise the emergent themes to construct robust findings.
The analysis mainly followed Charmaz’s constructivist grounded theory approach (Charmaz, 2006) and used Atlas.ti 111https://atlasti.com/. The first author began with line-by-line open coding on the interview transcripts, followed by focused coding to synthesize and organize the initial open codes into categories. He compiled analytical memos after each coding phase to document emerging categories, reflect on assumptions, and enable constant comparisons across participants, codes, and categories. The process was iterative and collaborative with other authors, leading to 101 codes in 9 categories. For example, the code VALUE: process within the VALUE category features discussions around artistic values inherent in creating and presenting robotic art. The code RELATIONSHIP: instrument within the RELATIONSHIP category illustrates how the artists perceive robots as instruments for their artistic pursuit. The numbers of codes and definitions of all categories are listed in Appendix A. Subsequently, we assigned high-level theoretical codes to the categories to delineate the relationships among the categories. This theoretical coding resulted in four themes explaining the generation of creative values in robotic art practice: “Embodied and material presence,” “Technological malfunction,” “Audience reception and feedback,” and “Creation and exhibition processes.” The iterative process forged us to construct a coherent narrative that well-represents the data. The analysis reasonably reached saturation as participants recruited later to a large extent repeated previous artists when answering questions regarding our key themes.
When the analysis was complete, we conducted a brief member check—shared our key findings (themes and explanations without quotations) with the artists, asked for comments, and revised the findings when deemed appropriate. Nevertheless, our positionalities inevitably shaped the analysis. The first and third authors have limited experience in professional art practices, which might hinder their ability to fully grasp the nuances of the artistic process. The second author, with an educational background in design and arts and publications in art exhibitions, brings a complementary artistic understanding to the analysis.
@@ 4. Findings
The inductive analysis across different robotic artists revealed recurrent factors that contribute to artistic creativity in robotic artwork. Here we present four such facets—Embodiment and Materiality, Malfunction, Audience’s Reaction and Reception, and Process of Creation and Exhibition. Robotic art is unique in each of them. We argue that these aspects are salient in the real-world practices of robotic art—uses and roles of robots in creative or creation activities. By investigating the practice of robotic art, our study contributes to an empirical understanding of how computing machines are creatively used for artistic and non-pragmatic purposes. Building upon prior works on artistic input to HCI (Kang et al., 2022), we advance the discourse by exploring how artistic practices, values, attitudes, and ways of thinking can serve as resources for HCI practitioners studying or designing for creative activities with machines.
@@ 4.1. Embodiment and Materiality
We found that the embodied nature of robotic art often becomes a source of inspiration for new artistic ideas. Linda, an artist-engineer who conducts scholarly research at the intersection of robotics and dance, reflected on how interacting with embodied robots makes her think about the differences between human and robotic bodies.
@@
The design of ChatGPT aligns with the mainstream approach to disembodied chatbots running as computer programs. In these instances, symbolic content (e.g., text, images, videos, audio) mostly serves as the communicative medium, as bodily interaction is minimized (i.e., users primarily sit and type). While creativity is arguably rooted in embodied interaction with other material bodies, current interactive agents (e.g., Copilot and Midjourney) designed to support creative work remain largely symbolic and disembodied. Limiting human-machine communication to symbolic channels may lose the benefits of embodiment in acquiring creativity.
@@
Embodiment has been an essential prerequisite of creativity for some artists since their creative production requires understanding embodied entities. To summarize this subsection, embodiment is such an important source of creativity for robotic artists. Practically, it yields new artistic expressions and aesthetics whose complexity is difficult to replicate by computer programs. The embodied form of robots, in turn, inspires creative ideas for artworks. These ideas can arise from understanding the entities embodied in the physical world, whether they are robots, humans, or other bodies in the environment. The symbolic and disembodied modes of interaction between human creators and machines in creative activities can be complemented and strengthened by embodied interaction.
@@ 4.2. Malfunction: “Ghost in the Machine”
Robots, encompassing both mechanical and computing devices, are inherently susceptible to malfunction, with physical robots being more prone to errors, glitches, and noise than virtual agents. These malfunctions are widespread in robotics. In robotic art, such errors hold unique significance, influencing the interpretation and value of the art. Unlike engineers, who aim to fix errors, robotic artists often embrace malfunctions as part of their creative process (N=7).
@@
Evelyn’s approach challenges the conventional expectation of machines as flawless and efficient executors. By slowing down the machine and celebrating its errors, she imbues the machine with a human-like quality of imperfection. This perspective turns the machine into something capable of growth and learning, much like a human child. The errors, therefore, potentially become a source of uniqueness and individuality in the artwork, adding depth and complexity to the artistic expression. This “counterintuitive” way of viewing error resonates with Alex who contrasts this view with the tendency of engineering that maintains neat, rational, and organized through monitoring and fixing errors: “[S]ometimes, it’s (error/glitch is) like a source of treasure. Like you find something that you could work on, you find something that people don’t really use… But when we are tinkering, we sometimes reach this point of, ‘ah, okay, now this is visible.’… I think sometimes even just those things (errors/glitches) could be a work of [art].”
@@
Many of our artists described how they perceive, evaluate, and appreciate the unexpectedness of robotic art, revealing new artistic ideas that would not have emerged otherwise: “[I]nstead of an ink particle, you had a hole in the form of that part… I was like, ‘Oh, we’ll see that the material is saturated, I will not push it.’ But the robot doesn’t have this understanding and pushes it. And I thought, ‘Oh, it’s actually a good outcome. It’s actually both conceptually and aesthetically very pleasing to me’” (Sophie).
In this case, the robot performed an action that a human artist normally would not perform. The robot made an unusual decision and breakthrough in expression, called by many artists as “surprise.” Once the artist recognizes its artistic value, it may be further explored and developed. Linda shared a similar anecdote where an unexpected jitter from the way the motor pulls a string gives a “texture” to the robotic movement, which she sees as creative.
@@
The fragility of robots is a widely shared concern among our robotic artists (N=8). Regardless of their origin—self-built, modified, or off-the-shelf—all robots are susceptible to breakdowns in real-world environments, particularly during extended exhibitions without proper maintenance by artists or qualified personnel. For exhibitions, malfunctions are generally unacceptable, and robots ought to perform flawlessly when showcasing to the audience. To address malfunctions, the artists came up with different strategies, such as having backup materials for replacement and assembling the robots on-site at the exhibition. One strategy is reducing the complexity of the robotic system, simplifying it to minimize the risk of failure or loss of control. Their approach involves designing robots that resist internal breakdowns and withstand external environmental factors, such as moisture and gravity. Linda explained, “If I do build them (robots), I try to keep them simple and I try to make something that will withstand its environment… Sometimes that might be outdoors next to the ocean for six days.” David further emphasized the need for simplicity in robotic design, stating,
@@ 4.3. Audience’s Reaction and Reception
The artistic and creative value in robotic artwork is determined not just by the work itself but often by the audience’s reactions and interpretations. Our artists (N=7) mentioned that they observe or think about audience reaction, and often incorporate them into subsequent iterations of their work. Alex, for instance, is motivated by seeing how people react to robots, drawing inspiration from their perceptions.
@@
People project something like animals or themselves or something [on the robots]. And then I got inspiration from that. Then I made a little brighter setup with some objects, a little bit like forest kind of setup. And then people try to imagine more stories. And then I also put some effect to [make the setting] looks like night or daylight or morning. Then people really see differently.
@@
Beyond individual interpretations, the way of interpretation is also socially shaped. Samuel made the point that the perception of creativity is also partly a social product because “creativity is depending on what kind of society we are in and what kind of people we are interacting with.” Mark and Robert extended that perception of robotic art is culturally conditioned, varying across different societies and generations. They mentioned how the animist cultural tendency of some East Asian societies potentially makes people more accepting and interested in robots and non-human entities (e.g., plants and animals) behaving as if are intelligent and agentic. The way that the social context of interpretation and perception determines artistic values reiterates our claim that the audience’s reception of robotic artwork is one of the key aspects of robotic art practice. It suggests that in achieving certain artistic goals by robotic art, considering the audience’s background and “horizon of expectations” (Jauss et al., 1982)—the socially and historically conditioned structure by which a person comprehends, interprets, and appraises any text based on cultural codes and lived experiences—may be constructive in refining the work’s idea.
@@ 4.4. Process of Creation and Exhibition
Many of the artists we interviewed (N=6) emphasized, or alluded to, the artistic value in the process of making robotic art. Specifically, there are two types of processes discussed here—the process of creation and the process of exhibition, reflecting two salient temporal stages of robotic art practice. We do not, by any means, suggest that process is unique to robotic art; apparently, other forms of art also attend to processes of their art practice. Our intention has been to situate the analysis of process in the emerging, particular context of robotic art and to reveal how process leads to a new understanding of robot’s uses and roles in real-world scenarios.
@@
This case illustrates how robotic artwork is not fixed but remains malleable even during the exhibition stage; temporal changes within the artwork can introduce new artistic qualities that evolve the work beyond its original design. Highlighting the artwork’s temporality here allows for elucidating how the current state of the created artifact and perceived creativity come to be. The practice of robotic art thus extends beyond the creation stage, encompassing the exhibition period. While in many cases the creation process is well planned, and temporal changes during the exhibition are typically unforeseen, both processes reveal that robotic art is in a state of ongoing creation across time. By paying attention to these processes, we unravel the temporal dimension that contributes to the creative values in robotic art.
In this Findings section, we have highlighted four aspects of robotic art practice that contribute to the artistic quality of the work or to achieving some artistic goals. The analysis reveals how various actors—artists, robots, audiences, and environments—are involved in the practice, influencing one another. These interactive patterns explain how creativity in robotic art is distributed within and emerges from the relations of actors. This idea echos with Daniel’s reflection, as he noted that he sees robotic artwork as “a product of a situation of a creative potential that is part of the environment, all the entities involved as well as me,” emphasizing the collaborative and emergent nature of creativity in robotic art.

### v2

@@ 3.1. Participant Recruitment
In this study, we conducted in-depth, semi-structured interviews with nine artists working on robotic art. They are renowned professional artists who use or build physical robots for their artistic work, namely robotic art. Designers and amateurs are out of this criterion. We identified potential participants by browsing individual and artistic organizations’ websites during the initial recruitment. Subsequently, we employed snowball sampling to recruit more artists who met our criteria. The recruitment ended when iterative data analysis reached saturation. The vast majority of the participants have had formal art training in drawing, painting, installation, media art, music, sculpture, or dance, and all of them incorporated robots into their particular genres. Some have technical backgrounds; many have a doctoral degree in the fields of art or technology. All of them build their robots or at least modify out-of-shell robots. Table 1 shows more demographic information. The study is approved by the Institutional Review Board of the lead author’s institution. The artists provided oral consent to participate in this study voluntarily.
@@ 3.2. Semi-Structured Interviews
Before the interviews, the first author reviewed the media press on the artists’ websites, gaining background information to build rapport, develop interview questions, and inform analysis. He conducted the interviews via Zoom with video recording, which lasted an average of 67 minutes. The in-depth interviews enabled us to look at the layers of meaning that artists relate with their work—unraveling feelings, attitudes, values, and motivations crucial for addressing our research question (Pugh, 2013). The initial interview questions were structured around the following topics: the workflow, motivation, and rewards and challenges of working on robotic art; the relationship between artists and robots; and the source of creativity. These inquiries were driven by our curiosity in making sense of the artists’ work practices. We aimed to understand their motivations for engaging with robotic art, unpacking the values embedded in their practices and informing broader topics around human-machine creation.
As the interviews began, the first author asked the artists to introduce themselves and their artwork. Then he asked about their inspiration source for their artwork and motivations for working on robotic art, steering the conversation toward creativity and the characteristics of robotic art. Following that, he solicited their reflections and anecdotes on the opportunities and challenges of working on robotic art to understand the nuances of robotic art practices. He then guided the conversation to explore the relationships between artists and robots, focusing on their respective agencies and contributions to creative outcomes in art practices. Additionally, he posed questions regarding the artistic practices’ embodied/material engagement. Concluding the session, he asked the artists to elaborate on any aspects they wished to add. To comply with the semi-structured nature, he prompted follow-up questions throughout the interviews and remained open to new ideas introduced by the artists, allowing some emergent ideas to lead the conversation. After each interview, the artist received a $40 Amazon or Visa e-gift card if they chose to accept compensation.
@@ 3.3. Qualitative Analysis
Data collection and analysis proceeded in parallel and iterative ways throughout the study. We began the analysis shortly after conducting each interview. Emerging insights from the analysis led us to refine our interview protocol several times, focusing on the most salient story in the data. The analysis also informed theoretical sampling (Charmaz, 2006) at the later stage, purposefully recruiting participants who are more likely to provide missing components or revise the emergent themes to construct robust findings.
The analysis mainly followed Charmaz’s constructivist grounded theory approach (Charmaz, 2006) and used Atlas.ti. The first author began with line-by-line open coding on the interview transcripts, followed by focused coding to synthesize and organize the initial open codes into categories. He compiled analytical memos after each coding phase to document emerging categories, reflect on assumptions, and enable constant comparisons across participants, codes, and categories. The process was iterative and collaborative with other authors, leading to 101 codes in 9 categories. For example, the code VALUE: process within the VALUE category marks discussions around artistic values inherent in creating and presenting robotic art. The code RELATIONSHIP: instrument within the RELATIONSHIP category illustrates how the artists perceive robots as instruments for their artistic pursuit. The numbers of codes and definitions of all categories are listed in Appendix A. Subsequently, we assigned high-level theoretical codes to the categories to delineate the relationships among the categories. This theoretical coding resulted in four themes explaining the generation of creative values in robotic art practice: Embodied and material presence, Technological malfunction, Audience reception and feedback, and Creation and exhibition processes. The iterative process forged us to construct a coherent narrative that well-represents the data. The analysis reasonably reached saturation as participants recruited later to a large extent repeated previous artists when answering questions regarding our key themes.
When the analysis was complete, we conducted a member check—shared the Findings section with the artists, asked for comments, and revised the findings when appropriate. Nevertheless, our positionalities inevitably shaped the analysis. The first and third authors have limited experience in professional art practices, which might constrain their ability to fully grasp the nuances of the artistic process. The second author, with an educational background in design and arts and publications in art exhibitions, brings complementary artistic expertise to the analysis.
@@ 4. Findings
The inductive analysis across different robotic artists revealed recurrent factors that contribute to artistic creativity in robotic artwork. Here we present four such facets—Embodiment and Materiality, Malfunction, Audience’s Reaction and Reception, and Process of Creation and Exhibition. Robotic art is unique in each of them. We argue that these factors are salient in the real-world practices of robotic art—uses of robots in artistic or creation activities. By investigating the practice of robotic art, our study contributes empirically to understanding how computing machines are creatively used for artistic and non-pragmatic purposes. Building upon prior works on artistic input to HCI (Kang et al., 2022), we advance the discourse by exploring how artistic practices, values, attitudes, and ways of thinking can serve as resources for HCI practitioners studying or designing for creative activities with machines.
@@ 4.1. Embodiment and Materiality
We found that the embodied nature of robotic art often becomes a source of inspiration for new artistic ideas. Linda, an artist-engineer who conducts scholarly research at the intersection of robotics and dance, reflected on how interacting with embodied robots makes her think about the differences between human and robotic bodies:
@@
The design of ChatGPT aligns with the mainstream approach to disembodied chatbots running as computer programs. In these designs, symbolic content (e.g., text, images, videos, audio) serves as the communicative medium, but bodily interaction is minimized (i.e., users primarily sit and type). While creativity is arguably rooted in embodied interaction with other material bodies, current interactive agents (e.g., Copilot and Midjourney) designed to support creative work remain largely symbolic and disembodied. Limiting human-machine communication to symbolic channels may lose the benefits of embodiment in acquiring creativity.
@@
Embodiment has been an essential prerequisite of creativity for some artists since their creative production requires understanding embodied entities. To summarize this subsection, embodiment is an important source of creativity for robotic artists. Practically, it yields new artistic expressions and aesthetics whose complexity is difficult to replicate by computer programs. The embodied form of robots, in turn, inspires creative ideas for artworks. These ideas can arise from understanding the entities embodied in the physical world, whether robots, humans, or other bodies in the environment. The symbolic and disembodied modes of interaction between human creators and machines in creative activities can be complemented and strengthened by embodied interaction.
@@ 4.2. Malfunction: “Ghost in the Machine”
Robots, encompassing both mechanical and digital devices, are inherently susceptible to malfunction, with physical robots being more prone to errors, glitches, and noise than virtual agents. These malfunctions are widespread in robotics. In robotic art, such errors hold unique significance, influencing the interpretation and value of the art. Unlike engineers, who aim to fix errors, robotic artists often embrace malfunctions as part of their creative process (N=7).
@@
Evelyn’s approach challenges the conventional expectation of machines as flawless and efficient executors. By slowing down the machine and celebrating its errors, she imbues the machine with a human-like quality of imperfection. This perspective turns the machine into something capable of growth and learning, much like a human child. The errors, therefore, potentially become a source of uniqueness and individuality in the artwork, adding depth and complexity to the artistic expression. This “counterintuitive” way of viewing error resonates with Alex who contrasts this view with the engineering tendency that strives to be neat, rational, and organized through monitoring and fixing errors: “[S]ometimes, it’s (error/glitch is) like a source of treasure. Like you find something that you could work on, you find something that people don’t really use… But when we are tinkering, we sometimes reach this point of, ‘ah, okay, now this is visible.’… I think sometimes even just those things (errors/glitches) could be a work of [art].”
@@
Many of our artists described how they perceive, evaluate, and appreciate the unexpectedness of robotic art, revealing new artistic ideas that would not have emerged otherwise: “[I]nstead of an ink particle, you had a hole in the form of that part… I was like, ‘Oh, we’ll see that the material is saturated, I will not push it (brush).’ But the robot doesn’t have this understanding and pushes it. And I thought, ‘Oh, it’s actually a good outcome. It’s actually both conceptually and aesthetically very pleasing to me’” (Sophie).
In this case, the robot performed an action that a human artist normally would not perform—pushing the brush on the canvas. The robot made an unusual decision and breakthrough in expression, called by many artists as “surprise.” Once the artist recognizes its artistic value, it may be further explored and developed. Linda shared a similar anecdote where an unexpected jitter from the way the motor pulls a string gives a “texture” to the robotic movement, which she sees as creative.
@@
The fragility of robots is a widely shared concern among our robotic artists (N=8). Regardless of their origin—self-built, modified, or off-the-shelf—all robots are susceptible to breakdowns in real-world environments, particularly during extended exhibitions without proper maintenance by artists or qualified personnel. For exhibitions, malfunctions are generally unacceptable, and robots ought to perform flawlessly when showcasing to the audience. To address malfunctions, the artists came up with different strategies, such as having backup materials for replacement and assembling the robots on-site at the exhibition. One strategy is reducing the complexity of the robotic system, simplifying it to minimize the risk of failure or loss of control. Their approach involves designing robots that resist internal breakdowns and withstand external environmental factors, such as moisture and gravity. Linda explained, “If I do build them (robots), I try to keep them simple and I try to make something that will withstand its environment… Sometimes that might be outdoors next to the ocean for six days.” David further emphasized that the concern for fragility leads to the need for simplicity in robotic design:
@@ 4.3. Audience’s Reaction and Reception
The artistic and creative value in robotic artwork is determined not just by the work itself but often by the audience’s reactions and interpretations. Our artists (N=7) mentioned that they observe or think about audience reaction, and often incorporate them into subsequent iterations of their work. Alex, for instance, is motivated in the first place by observing how people react to robots, drawing inspiration from their perceptions.
@@
People project something like animals or themselves or something [on the robots]. And then I got inspiration from that. Then I made a little brighter setup with some objects, a little bit like forest kind of setup. And then people try to imagine more stories. And then I also put some effect to [make the setting] looks like night or daylight or morning. Then people really see [the robots] differently.
@@
Beyond the individual level, the way of interpretation is also socially shaped. Samuel made the point that the perception of creativity is also partly a social product because “creativity is depending on what kind of society we are in and what kind of people we are interacting with.” Mark and Robert extended that the perception of robotic art is culturally conditioned, varying across different societies and generations. They mentioned how the animist cultural tendency of some East Asian societies potentially makes people more willing to accept and interested in robots and non-human entities (e.g., plants and animals) behaving as if intelligent and agentic. The way that the social context of interpretation and perception determines artistic values reiterates our claim that the audience’s reception of robotic artwork is one of the key aspects of robotic art practice. It suggests that in achieving certain artistic goals by robotic art, considering the audience’s background and “horizon of expectations” (Jauss et al., 1982)—the socially and historically conditioned structure by which a person comprehends, interprets, and appraises any text based on cultural codes and lived experiences—may be constructive in refining the work’s idea.
@@ 4.4. Process of Creation and Exhibition
Many of the artists we interviewed (N=6) emphasized, or alluded to, the artistic value in the process of making robotic art. Specifically, two types of processes are discussed here—the process of creation and the process of exhibition, reflecting two salient temporal stages of robotic art practice. We do not, by any means, suggest that process is unique to robotic art; apparently, other forms of art also attend to processes of their art practice. Our intention has been to situate the analysis of process in the emerging, particular context of robotic art and to reveal how process leads to a new understanding of robot’s uses and roles in real-world scenarios.
@@
This case illustrates how robotic artwork is not fixed but remains malleable even during the exhibition stage; temporal changes within the artwork can introduce new artistic qualities that evolve the work beyond its original design. Highlighting the artwork’s temporality here allows for elucidating how the current state of the created artifact and creativity come to be. The practice of robotic art thus extends beyond the creation stage, encompassing the exhibition period. While in many cases the creation process is well planned, and temporal changes during the exhibition are typically unforeseen, both processes reveal that robotic art is in a state of ongoing creation across time. By paying attention to these processes, we unravel the temporal dimension that contributes to the creative values in robotic art.
In this Findings section, we have highlighted four aspects of robotic art practice that contribute to the artistic quality of the work or to achieving some artistic goals. The analysis reveals how various actors—artists, robots, audiences, and environments—are involved in the practice, influencing one another. These interactive patterns explain how creativity in robotic art is distributed within and emerges from the relations of actors. This idea echos with Daniel’s reflection, as he noted that he sees robotic artwork as “a product of a situation of a creative potential that is part of the environment, all the entities involved as well as me,” emphasizing the distributed and emergent nature of creativity in robotic art.